# How to Synthesize Text Data without Model Collapse?

### v1

@@ Abstract
Model collapse in synthetic data indicates that iterative training on self-generated data leads to a gradual decline in performance. With the proliferation of AI models, synthetic data will fundamentally reshape the web data ecosystem. Future GPT-$\{n\}$ models will inevitably be trained on a blend of synthetic and human-produced data. In this paper, we focus on two questions: what is the impact of synthetic data on language model training, and how to synthesize data without model collapse? We first pre-train language models across different proportions of synthetic data, revealing a negative correlation between the proportion of synthetic data and model performance. We further conduct statistical analysis on synthetic data to uncover distributional shift phenomenon and over-concentration of n-gram features. Inspired by the above findings, we propose token editing on human-produced data to obtain semi-synthetic data. As a proof of concept, we theoretically demonstrate that token-level editing can prevent model collapse, as the test error is constrained by a finite upper bound. We conduct extensive experiments on pre-training from scratch, continual pre-training, and supervised fine-tuning. The results validate our theoretical proof that token-level editing improves data quality and enhances model performance.
Figure 1: Model collapse of synthetic data.  The model continuously trains on its previously generated data, leading to a gradual decline in model performance, i.e., model collapse. Starting from real data $(x_{o},y_{o})$, the test error $E_{test}$ increases as $f_{0}$ undergoes iterative training on synthetic data $(y_{1},y_{2},\dots,y_{n})$. ToEdit (ours), we use a trained model for token-level editing rather than purely synthesizing data. Leveraging $f_{0}$ and an operation matrix $m_{i}$ to edit the data, the test error is constrained within a fixed upper bound. Therefore, we can preserve the distribution coverage to avoid model collapse.
@@ 1 Introduction
Model collapse refers to a degenerative process in which the output data of learned generative models contaminates the training sets of subsequent generations. As shown in Figure 1, iterative training coupled with data synthesis induces a progressive accumulation of test errors (Shumailov et al., 2024; Dohmatob et al., 2024a). Consequently, generative models increasingly overfit to synthetic data distributions, failing to capture the complexity in human-produced data. Through successive iterations in Figure 1, these distortions accumulate, finally undermining the model’s capacity for generalization.
@@
Recent studies focus on two aspects. First, theoretical foundations of model collapse. Shumailov et al. (2024) and Dohmatob et al. (2024a) identify the model collapse phenomenon and formalize a theoretical framework based on linear regression. Gerstgrasser et al. (2024) demonstrate that if synthetic data is accumulated while retaining the initial real data, the test error will be bounded, thus breaking model collapse. Dohmatob et al. (2024c; b) indicate that missing long tails of synthetic data lead to scaling law cutoff. Second, practical implementations on synthetic datasets by diverse prompting. Synthetic datasets (Trinh et al., 2024; Cui et al., 2023; Zhang et al., 2024) have been proven to boost the capabilities of language models. Cheng et al. (2024a; b); Maini et al. (2024) rephrase text into more formal styles, thereby improving the data quality. There are still two key questions that require further investigation:
(Q1) What is the impact of synthetic data on language model training?
(Q2) How can data be synthesized without causing model collapse?
@@
In this paper, we address the first question by training language models on varying mixtures of synthetic and human-produced data, demonstrating non-iterative model collapse. Unlike the original model collapse setting which iteratively trains on self-generated data, we directly mix synthetic and human-produced data to create training datasets with different mixing ratios. The results show a negative correlation between performance and the proportion of synthetic data. Subsequent statistical analysis on distributions and features indicates coverage collapse—synthetic data covers only a small portion of the human-produced data distribution—and over-concentration of synthetic n-gram features. Based on the above findings, we address the second question by proposing token-level editing (ToEdit), resamples and replaces data points with relatively high model confidence. As illustrated in Figure 1, ToEdit preserves distribution coverage and theoretically constrains test error within a fixed upper bound. Furthermore, extensive experiments across pre-training from scratch, continual pre-training, and supervised fine-tuning confirm its positive impact on model performance.
@@
Contributions. We summarize the key contributions of this work as follows:
  • We demonstrate non-iterative model collapse by pre-training language models on a mixture of synthetic and human-produced data (§ 2.1): directly mixing pure synthetic data [^1] In short, pure synthetic data is generated entirely by language models, while semi-synthetic data comes from limited modifications to human-produced data. Detailed definitions are provided in § 5.2., without iterative training, leads to performance degradation.
@@
  • We perform a distributional statistical analysis, revealing that synthetic data leads to coverage collapse and over-concentration of n-gram features. Even subsequent data selection struggles to correct the distribution (§ 2.2).
@@
Table 1: PPL evaluation results for GPT-2 Small (124M) pre-trained on data mixture. The PPL increases as the proportion of synthetic data grows, providing further confirmation of Figure 2.
ArXiv Books2 Books3 Math Enron EuroParl FreeLaw GitHub PG-19 HackerNews NIH
Human data 22.26 25.39 22.87 10.84 23.50 30.73 12.04 4.15 16.88 32.54 23.53
25% Synthetic Data 21.86 26.32 23.87 11.05 24.85 35.02 12.84 4.35 17.99 33.80 23.76
50% Synthetic Data 22.50 28.01 25.75 10.84 26.56 41.99 14.02 4.67 19.70 36.12 24.61
75% Synthetic Data 24.35 31.19 28.98 11.81 30.30 56.32 16.03 5.30 22.75 40.44 26.19
Synthetic Data 35.60 43.72 47.72 17.25 66.97 129.75 29.62 12.00 50.14 87.95 39.48
@@
Human data 28.08 25.77 33.56 26.78 18.97 15.49 10.81 20.86 19.32 24.31 21.54 21.37
25% Synthetic Data 29.25 26.94 34.63 27.83 19.55 15.38 11.03 22.32 19.58 25.88 22.63 22.31
50% Synthetic Data 31.00 28.76 37.48 29.36 20.51 15.89 11.54 23.53 20.51 27.57 24.91 23.90
75% Synthetic Data 34.18 32.04 42.39 32.17 22.33 16.92 12.55 26.54 22.21 30.68 28.98 27.03
Synthetic Data 57.83 53.94 78.18 54.69 34.82 23.87 20.47 51.78 37.24 46.12 65.49 49.30
Figure 2: Non-iterative model collapse. Training language models from scratch on AI-synthesized data or a mixture of human and synthetic data leads to performance degradation. This degradation is negatively correlated with the proportion of synthetic data used in training. A. We pre-train GPT-2 Small (124M) on human (Dolma (Soldaini et al., 2024)) and synthetic (Cosmopedia (Ben Allal et al., 2024)) data. As the proportion of synthetic data increases, the model’s loss decreases. B. As the proportion of synthetic data increases, the PPL also rises. This trend remains consistent across different validation sets. More results on downstream tasks are presented in 10 and  11.
@@ 2 Non-iterative Model Collapse
In this section, we conduct pre-training on synthetic data mixtures and explore the reasons behind non-iterative model collapse. Prior studies (Shumailov et al., 2024; Dohmatob et al., 2024a) investigate the curse of recursion, where iterative training on self-generated data leads to a degenerative process known as iterative model collapse. In contrast, we investigate non-iterative model collapse, training a model directly on data synthesized by other models. More discussion is provided in § 5.1.
@@ 2.1 Pre-training on Data Mixture
Extensive prior work have proved synthetic data can boost language models’ capabilities, including instruction following (Wang et al., 2022), reasoning (Zhu et al., 2023; Trinh et al., 2024), alignment (Cui et al., 2023), biomedicine (Zhang et al., 2024) and so on. In this section, we investigate the impact of synthetic data on pre-training. Compared with prior studies focused on SFT and RLHF, we examine synthetic data integration in a more fundamental stage of language model.
Setup We define the mixing ratio between human-produced and synthetic data as $\alpha$, where $0\leq\alpha\leq 1$. The total amount of training data $D_{\text{total}}$ is expressed as a combination of human-produced data $D_{\text{human}}$ and synthetic data $D_{\text{synthetic}}$, represented by the formula: $D_{\text{total}}=\alpha D_{\text{human}}+(1-\alpha)D_{\text{synthetic}}$.
@@
We use Dolma (Soldaini et al., 2024) as source human-produced data. We use Cosmopedia (Ben Allal et al., 2024) as the source synthetic data, which is distilled from Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). We train GPT-2 (Radford et al., 2019) and OLMo (Groeneveld et al., 2024) from scratch, using data mixtures containing 50B tokens each. The PPL evaluation sets are drawn from Paloma benchmark (Magnusson et al., 2023) [^2] Paloma-bench is specifically designed for PPL evaluation, particularly for Dolma. And the Dolma team has already performed data de-contamination to prevent data leakage. and the 22 sub-domain validation sets of the Pile (Gao et al., 2020b), as used in (Maini et al., 2024).
@@
Finding I: Incorporating pure synthetic data harms the language models pre-training. Main results are presented in Figure 2, the PPL on real-world validation sets increases as the proportion of synthetic data grows, indicating degraded model performance. When training from scratch, synthetic data does not benefit the model and may even hinder its learning process. However, incorporating human-produced data into the training mixture mitigates model collapse to some extent. Compared to previous research on iterative model collapse (Shumailov et al., 2024; Dohmatob et al., 2024a; c), the non-iterative damage caused by synthetic data is more concerning and directly relevant to the training of next-generation language models.
@@
Figure 3: PPL distribution of human and synthetic data estimated by Llama-3-8B. The synthetic data lacks the long tail of the human-produced data and is also concentrated within the first $25\%$ of the human-produced data distribution. A. Distribution of human-produced data is sharp with a long tail, spanning a wide range from 0 to over 100. B. The values are concentrated within a much narrower range, mostly between 0 and 12. The experiment uses Dolma v6 and Cosmopedia as human and synthetic data, each with sampled 6B tokens. More results in Figure 9.
2.2 Why does synthetic data fail in language model pre-training?
@@
We conduct three statistical analyses: (1) sample-level distribution, (2) feature-based overlap, and (3) distribution-reference data selection. The experimental results reveal that, compared to human-produced data, synthetic data lacks long-tail coverage and suffers from coverage collapse. The limited diversity and concentrated features in synthetic data make using human-produced data as a reference to select synthetic data particularly challenging.
@@
Setup We conduct statistical and feature-based analyses to explore why synthetic data fails in pre-training. (1) We leverage a prior distribution to estimate the human-produced and synthetic data. We use Llama-3-8B (AI@Meta, 2024) and StableLM-Zephyr-3B (Bellagente et al., 2024). Different priors consistently yield the same results. (2) We analyze the n-gram features of human-produced and synthetic data from a feature-based perspective, such as n-gram response values. (3) Based on the distribution of human-produced data, we sample data from the synthetic dataset that closely matches the human-produced data distribution in an attempt to filter the synthetic data.
@@
Finding II: Synthetic data distribution not only misses long tails, but also causes coverage collapse. Figure 3 and 9 illustrate that the PPL of synthetic data is confined to the lower $25\%$ of the human-produced data, failing to capture the full range and complexity of human-produced data distributions. Specifically, as illustrated in Figure 3A, human-produced data exhibit a wide distribution in the range $[1,100+]$, characterized by a sharp peak and a pronounced long tail. In contrast, as shown in Figure 3B, the synthetic data is confined to a narrower range of $[0,14]$, displaying a smoother distribution. Additional results of StabLM are shown in Figure 9. While the absolute PPL ranges estimated by different models may vary, the relative shapes and proportional ranges of these two distributions remain consistent. This phenomenon provides evidence that when scaling up to larger synthetic datasets, there will be a notable absence of the long tail. In summary, we observe a more severe coverage collapse. This limited coverage reduces the data’s ability to generalize effectively and may contribute to model collapse, as shown in Figure 2.
@@
Figure 4: A. Embedding visualization using t-SNE and sentence-transformers. B. pre-training results for selected synthetic data and other data mixtures.
@@
Finding III: Synthetic data over-concentrates N-gram features. Based on the above distribution estimate, we further analyze why synthetic data fails at the feature level. Figure 10 and 11 demonstrate that synthetic data exhibits higher frequencies of certain bi-grams compared to human-produced data. To further examine feature-level differences, we hash unigram and bi-gram features into 10,000 hash buckets. As illustrated in Figure 5, human-produced data displays a noticeably broader response range, while synthetic data features are concentrated in a few specific buckets. This indirectly supports our earlier observation of feature over-concentration. We then expanded the hash bucket range to 1,000 × 20,000 buckets and used a locality-sensitive hashing method to differentiate the features more precisely. Nevertheless, the results remained consistent. As shown in Figure 12, the majority of the response values are close to zero. Distinguishing features in synthetic data remains challenging.

### v2

@@ Abstract
Model collapse in synthetic data indicates that iterative training on self-generated data leads to a gradual decline in performance. With the proliferation of AI models, synthetic data will fundamentally reshape the web data ecosystem. Future GPT-$\{n\}$ models will inevitably be trained on a blend of synthetic and human-produced data. In this paper, we focus on two questions: what is the impact of synthetic data on language model training, and how to synthesize data without model collapse? We first pre-train language models across different proportions of synthetic data, revealing a negative correlation between the proportion of synthetic data and model performance. We further conduct statistical analysis on synthetic data to uncover distributional shift phenomenon and over-concentration of n-gram features. Inspired by the above findings, we propose token editing on human-produced data to obtain semi-synthetic data. As a proof of concept, we theoretically demonstrate that token-level editing can prevent model collapse, as the test error is constrained by a finite upper bound. We conduct extensive experiments on pre-training from scratch, continual pre-training, and supervised fine-tuning. The results validate our theoretical proof that token-level editing improves model performance.
Machine Learning, ICML
Figure 1: Model collapse of synthetic data.  The model continuously trains on its previously generated data, leading to a gradual decline in model performance, i.e., model collapse. Starting from real data $Data_{0}$, the test error $E_{test}$ increases as $f_{0}$ undergoes iterative training on synthetic data $Data_{>0}$. ToEdit (ours), we use a trained model for token-level editing rather than purely synthesizing data. Leveraging $f_{0}$ and an operation matrix $M_{i}$ to edit the data, the test error is constrained within a fixed upper bound. Therefore, we can preserve the distribution coverage to avoid model collapse.
@@ 1 Introduction
Model collapse refers to a degenerative process in which the output data of learned generative models contaminates the training sets of subsequent generations. As shown in Figure 1, iterative training coupled with data synthesis induces a progressive accumulation of test errors (Shumailov et al., 2024; Dohmatob et al., 2024a). Consequently, generative models increasingly overfit to synthetic data distributions, failing to capture the complexity in human-produced data. Through successive iterations in Figure 1, these distortions accumulate, finally undermining the model’s capacity.
@@
Recent studies focus on two aspects. First, theoretical foundations of model collapse. Shumailov et al. (2024) and Dohmatob et al. (2024a) identify the model collapse phenomenon and formalize a theoretical framework based on linear regression. Gerstgrasser et al. (2024) demonstrate that if synthetic data is accumulated while retaining the initial real data, the test error will be bounded, thus breaking model collapse. Dohmatob et al. (2024c, b) indicate that missing long tails of synthetic data lead to scaling law cutoff. Second, practical implementations on synthetic datasets by diverse prompting. Synthetic datasets (Trinh et al., 2024; Zhang et al., 2024) have been proven to boost the capabilities of language models. Cheng et al. (2024a, b); Maini et al. (2024) rephrase text into more formal styles, thereby improving the data quality. There are still two key questions that require further investigation: (Q1) What is the impact of synthetic data on language model training? (Q2) How can data be synthesized without causing model collapse?
@@
In this paper, we address the first question by training language models on varying mixtures of synthetic and human-produced data, demonstrating non-iterative model collapse. Unlike the original model collapse setting which iteratively trains on self-generated data, we directly mix synthetic and human-produced data to create training datasets with different mixing ratios. The results show a negative correlation between performance and the proportion of synthetic data. Subsequent statistical analysis on distributions and features indicates coverage narrowing—synthetic data covers only a small portion of the human-produced data distribution—and over-concentration of synthetic n-gram features. Based on the above findings, we address the second question by proposing token-level editing (ToEdit), resamples and replaces data points with relatively high model confidence. As illustrated in Figure 1, ToEdit preserves distribution coverage and theoretically constrains test error within a fixed upper bound. Extensive experiments across pre-training from scratch, continual pre-training, and supervised fine-tuning confirm its positive impact on model performance.
@@
Contributions. We summarize the key contributions of this work as follows [^1] Code repository available at https://github.com/Xuekai-Zhu/toedit.:
@@
  • We demonstrate non-iterative model collapse by pre-training language models on a mixture of synthetic and human-produced data (§ 3.1): directly mixing pure synthetic data, without iterative training, leads to performance degradation. Furthermore, we perform a distributional statistical analysis, revealing that synthetic data leads to coverage narrowing and over-concentration of n-gram features. Even subsequent data selection struggles to correct the distribution (§ 3.2).
@@
2 Background
Shumailov et al. (2024); Dohmatob et al. (2024a, c) demonstrate AI models trained recursively on data generated by earlier versions of themselves can result in performance degradation, ultimately rendering the AI model completely useless. This process can be formulated as follows:
$E_{test}(\hat{w}_{n+1})=\frac{\sigma^{2}d}{T-d-1}\times n.$
This indicates that the error will continuously increase with the number of iterations $n$. The detailed theoretical notation is provided in § 4.2. Dohmatob et al. (2024c) further point out that synthetic data also contribute to a truncation of the scaling law. Gerstgrasser et al. (2024); Seddik et al. (2024) further adjust the data iteration setting to data accumulation or real data mixing. They demonstrate that data accumulation can prevent model collapse. Inspired by the above work, we further explore the impact of synthetic data in pre-training and analyze its differences from real data. Building on our findings, we propose token editing as a method to prevent model collapse during data synthesis. Further comparisons are in Appendix B and D.
Table 1: Subdomain PPL evaluation results for GPT-2 Small (124M) pre-trained on data mixture. The PPL increases as the proportion of synthetic data grows, providing further confirmation of Figure 2.
ArXiv Books2 Books3 Math Enron EuroParl FreeLaw GitHub PG-19 HackerNews NIH Avg
Human data 22.26 25.39 22.87 10.84 23.50 30.73 12.04 4.15 16.88 32.54 23.53 20.99
25% Synthetic Data 21.86 26.32 23.87 11.05 24.85 35.02 12.84 4.35 17.99 33.80 23.76 22.06
50% Synthetic Data 22.50 28.01 25.75 10.84 26.56 41.99 14.02 4.67 19.70 36.12 24.61 23.48
75% Synthetic Data 24.35 31.19 28.98 11.81 30.30 56.32 16.03 5.30 22.75 40.44 26.19 27.60
Synthetic Data 35.60 43.72 47.72 17.25 66.97 129.75 29.62 12.00 50.14 87.95 39.48 51.93
@@ 2 Background
Human data 28.08 25.77 33.56 26.78 18.97 15.49 10.81 20.86 19.32 24.31 21.54 22.59
25% Synthetic Data 29.25 26.94 34.63 27.83 19.55 15.38 11.03 22.32 19.58 25.88 22.63 23.91
50% Synthetic Data 31.00 28.76 37.48 29.36 20.51 15.89 11.54 23.53 20.51 27.57 24.91 25.09
75% Synthetic Data 34.18 32.04 42.39 32.17 22.33 16.92 12.55 26.54 22.21 30.68 28.98 28.64
Synthetic Data 57.83 53.94 78.18 54.69 34.82 23.87 20.47 51.78 37.24 46.12 65.49 47.87
Figure 2: Non-iterative model collapse. Training language models from scratch on AI-synthesized data or a mixture of human and synthetic data leads to performance degradation. This degradation is negatively correlated with the proportion of synthetic data used in training. Setting: We pre-train GPT-2 Small (124M) on human data (Dolma (Soldaini et al., 2024)) and synthetic data (Cosmopedia (Ben Allal et al., 2024)) and evaluate the PPL on the Paloma benchmark (Magnusson et al., 2023). Training loss in Figure 7. Further validations on 22 subdomains and general downstream tasks are presented in Table 1 and Table 9, respectively.
@@ 3 Non-Iterative Model Collapse
Prior studies (Shumailov et al., 2024; Dohmatob et al., 2024a) investigate the curse of recursion, where iterative training on self-generated data leads to a degenerative process known as iterative model collapse. However, we often face direct data mixing of human-produced and synthetic data, and pre-training from scratch. We attempt to analyze model collapse in this more general scenario, called non-iterative model collapse. Specifically, we conduct pre-training on synthetic data mixtures and explore the reasons behind non-iterative model collapse through data distribution and characteristics. A more detailed comparison of iterative and non-iterative settings is in Appendix G.1.
@@ 3.1 Pre-training on Data Mixture
In this section, we investigate the impact of synthetic data on pre-training. Compared with studies on SFT and RLHF, we examine synthetic data integration in a more fundamental stage of the language model.
@@
Setup We pre-train GPT-2 (Radford et al., 2019) and OLMo (Groeneveld et al., 2024) from scratch, using data mixtures containing 50B tokens each. We define the mixing ratio between human-produced and synthetic data as $\alpha$, where $0\leq\alpha\leq 1$. The total amount of training data $D_{\text{total}}$ is a combination of human-produced data $D_{\text{human}}$ and synthetic data $D_{\text{synthetic}}$, represented by the formula: $D_{\text{total}}=\alpha D_{\text{human}}+(1-\alpha)D_{\text{synthetic}}$. We use Dolma (Soldaini et al., 2024) as source human-produced data. We use Cosmopedia (Ben Allal et al., 2024) as the source synthetic data, which is distilled from Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). To ensure rigorous validation and prevent data leakage, we construct a three-tier evaluation: (1) the Paloma benchmark (Magnusson et al., 2023), including carefully decontaminated test sets for Dolma; (2) comprehensive PPL evaluation across 22 subdomains from the Pile (Gao et al., 2020b); and (3) seven general downstream tasks as outlined in (Maini et al., 2024).
@@
Finding I: Incorporating synthetic data harms the language models pre-training. PPL results of Paloma benchmark and 22 subdomains are presented in Figure 2 and Table 1, respectively. These results demonstrate that PPL on real-world validation sets increases as the proportion of synthetic data grows, indicating degraded model performance. When training from scratch, synthetic data does not benefit the model and may even hinder its learning process. However, incorporating human-produced data into the training mixture mitigates model collapse to some extent. Further results on general downstream tasks in Table 9 and 11 also corroborate the above findings. The overall trend shows a decline as the proportion of synthetic data increases, with models trained on purely synthetic data performing the worst. Compared to previous research on iterative model collapse (Shumailov et al., 2024; Dohmatob et al., 2024a, c), the non-iterative damage caused by synthetic data is more concerning and directly relevant to the training of next-generation language models.
@@
Figure 3: PPL distribution of human and synthetic data estimated by Llama-3-8B. The synthetic data lacks the long tail of the human-produced data and is also concentrated within the first $25\%$ of the human-produced data distribution. A. Distribution of human-produced data is sharp with a long tail, spanning a wide range from 0 to over 100. B. The values are concentrated within a much narrower range, mostly between 0 and 14. The same trend estimated by StableLM-3B is demonstrated in Figure 10.
3.2 Why Does Synthetic Data Fail in Pre-training?
@@
We conduct three statistical analyses: (1) sample-level distribution, (2) feature-based overlap, and (3) distribution-reference data selection. The experimental results reveal that, compared to human-produced data, synthetic data lacks the long-tail samples and suffers from coverage narrowing. The limited diversity and concentrated features in synthetic data make using human-produced data as a reference to select synthetic data particularly challenging.
@@
Setup We conduct statistical and feature-based analyses to explore why synthetic data fails in pre-training. (1) We leverage a prior distribution to estimate the human-produced and synthetic data. We use Llama-3-8B (AI@Meta, 2024) and StableLM-Zephyr-3B (Bellagente et al., 2024). Different priors consistently yield the same results. (2) We analyze the n-gram features of human-produced and synthetic data from a feature-based perspective, such as n-gram response values. (3) Based on the features of human-produced data, we apply importance sampling (Xie et al., 2023) to filter synthetic data that closely aligns with human-produced features. More details of importance sampling are in § G.5.
@@
Finding II: Synthetic data distribution not only lacks long tails but also exhibits significant coverage narrowing. Figure 3 illustrate that the PPL of synthetic data is confined to the lower $25\%$ of the human-produced data, failing to capture the full range and complexity of human-produced data distributions. Specifically, as illustrated in Figure 3A, human-produced data exhibit a wide distribution in the range $[1,100+]$, characterized by a sharp peak and a pronounced long tail. In contrast, as shown in Figure 3B, the synthetic data is confined to a narrower range of $[0,14]$, displaying a smoother distribution. Additional results of StabLM are shown in Figure 10. While the absolute PPL ranges estimated by different models may vary, the relative shapes and proportional ranges of these two distributions remain consistent. This phenomenon demonstrate that when scaling up to larger synthetic datasets, there will be a notable absence of the long tail, leading to severe coverage narrowing. This limited coverage reduces the generalization ability and contribute to model collapse.
@@
Figure 4: A. Pre-training results for selected synthetic data and other data mixtures on OLMo-237M. B. Embedding visualization between human-produced, synthetic, and DSIR-selected data using sentence-transformer.
@@
Finding III: Synthetic data over-concentrates N-gram features. Based on the above distribution estimate, we further analyze why synthetic data fails at the feature level. Figure 12 and 13 demonstrate that synthetic data exhibits higher frequencies of certain bi-grams compared to human-produced data. To further examine feature-level differences, we hash uni-gram and bi-gram features into 10,000 hash buckets. As illustrated in Figure 11, human-produced data displays a noticeably broader response range, while synthetic data features are concentrated in a few specific buckets. This indirectly supports our earlier observation of feature over-concentration. We then expanded the hash bucket range to 1,000 × 20,000 buckets and used a locality-sensitive hashing method to differentiate the features more precisely. The results remain consistent. As shown in Figure 14, most response values are near zero. Distinguishing features in synthetic data remains challenging.
Finding IV: Distribution shifting cannot be mitigated through data selection. Inspired by recent data selection works (Xie et al., 2023; Albalak et al., 2024), we try to leverage the human-produced data features as a reference distribution for selecting synthetic samples. We apply importance sampling from DSIR (Xie et al., 2023) to filter synthetic data. As illustrated in Figure 4A, the training results of selected synthetic samples still fluctuates around the original performance of the synthetic data, indicating that even biased sampling cannot correct the distributional shift. As shown in Figure 4B, the sampled data still fails to align with human-produced data in the embedding space, even at the boundary regions of the synthetic data.
3.3 Proposed Strategy
Following these lessons so far, due to the coverage and feature over-concentration properties of synthetic data, the best approach is to rely entirely on human-produced data and avoid including synthetic data. However, we are still wondering how synthetic data can be used to enhance human-produced data. This leads to a general guideline for synthetic data: relying solely on synthetic data leads to model collapse, so preserving the primary human-produced data distribution is essential. In that case, we propose token-level editing, which leverages a prior distribution to adjust the data. Our method can maintain the source distribution while improving the source data, called semi-synthetic data.